OpenStudy (moose213):

Please help will fan and medal! George is building a fence around a rectangular dog run. He is using his house as one side of the run. The area of the dog run will be 240 square feet. The length of the run is 30 feet, and the width is (30 - x) feet. Recall the formulas for area and perimeter: A = lw and P = 2l + 2w. How many feet of fencing will George need for the dog run? 22 feet 68 feet 76 feet 82 feet

OpenStudy (moose213):

@Qwertty123 @jjcoolja12 @Ms-Brains @563blackghost

OpenStudy (moose213):

Please help someone this is the last question that I need help on, the homework's due tomorrow pleaseee

563blackghost (563blackghost):

So we know that the area of the dog run is 240 square feet and that the length is 30 feet but the width is the length of the dog run minus x so lets set this up.... \(\Huge{240=30(30-x)}\) Lets simplify to find out what x equals.

OpenStudy (qwertty123):

Can you plug in the information that you have?

OpenStudy (jjcoolja12):

id say 68

OpenStudy (qwertty123):

lol nvm @563blackghost got it!

OpenStudy (moose213):

I think @jjcoolja12 is right 68..?

OpenStudy (qwertty123):

You need to simplify it. Here: Simplifying 240 = 30(30 + -1x) 240 = (30 * 30 + -1x * 30) 240 = (900 + -30x) Solving: 240 = 900 + -30x Move all terms containing x to the left, all other terms to the right. Add '30x' to each side of the equation. 240 + 30x = 900 + -30x + 30x

OpenStudy (jjcoolja12):

Let's solve your equation step-by-step. 240=30(30−1x) Step 1: Simplify both sides of the equation. 240=30(30−1x) Simplify: (Hide steps) 240=(30)(30)+(30)(−1x)(Distribute) 240=900+−30x 240=−30x+900 Step 2: Flip the equation. −30x+900=240 Step 3: Subtract 900 from both sides. −30x+900−900=240−900 −30x=−660 Step 4: Divide both sides by -30. −30x −30 = −660 −30 x=22 Answer: x=22

OpenStudy (jjcoolja12):

i was wrong it was 22

OpenStudy (jjcoolja12):

http://www.mathpapa.com/algebra-calculator.html

OpenStudy (qwertty123):

Then you Combine like terms: -30x + 30x = 0 240 + 30x = 900 + 0 240 + 30x = 900 Add '-240' to each side of the equation. 240 + -240 + 30x = 900 + -240 Combine like terms: 240 + -240 = 0 0 + 30x = 900 + -240 30x = 900 + -240

OpenStudy (moose213):

So 22 or 68 lmao

OpenStudy (qwertty123):

Combine like terms: 900 + -240 = 660 30x = 660 Divide each side by '30'. x = ?

OpenStudy (jjcoolja12):

?

OpenStudy (moose213):

@563blackghost 68?

OpenStudy (qwertty123):

Wait.. the question is asking...How many feet of fencing will George need for the dog run?

OpenStudy (moose213):

@Qwertty123 22

OpenStudy (qwertty123):

Yes 22! ^.^

OpenStudy (qwertty123):

I mean 22 feet! XD

OpenStudy (moose213):

Yea, it's asking How many feet of fencing will George need for the dog run?

OpenStudy (moose213):

Thanks fam!

OpenStudy (qwertty123):

So you got the answer :)

OpenStudy (moose213):

Yea thanks!

OpenStudy (moose213):

Don't worry, I got you on the medal.

OpenStudy (qwertty123):

\(\color{orange}{Your~Welcome!}\)

OpenStudy (moose213):

And fanned.

OpenStudy (qwertty123):

Thanks <3

OpenStudy (moose213):

Alright cya!